OpenStudy (waheguru):

How Do I solve the problem A diver dives into a pool. His depth is given by the equation d = 0.5(t – 1)(t – 20), where d is his depth relative to ground level, and t is the time in seconds after he entered the water. a) What is the length of time the driver travels before he resurfaces?

OpenStudy (freethinker):

can you try?

OpenStudy (waheguru):

I did but Im not sure

OpenStudy (waheguru):

We did this as an example in class and I wanted to ogver it

OpenStudy (waheguru):

The time he traveled is 19seconds but I dont understand why?

OpenStudy (freethinker):

did you copy the solution?

OpenStudy (waheguru):

0=0.5(t-1)(t-20) t - 1 = 0 t = 1 and t - 20 = 0 t = 20

OpenStudy (waheguru):

Then we subtracted 20 - 1 which I did not understand

OpenStudy (freethinker):

why didn't you ask then?